# aoas/sm_enkf

### v2

@@ 3.1 Consistent estimation of high dimensional $\boldsymbol{\Sigma}_{t}^{f}$
    For the mid-banding estimator $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$ with $k_{1}+k_{2}\asymp(n^{-1}\log p)^{-1/(2(\alpha+1))}$, then uniformly on $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
@@
    For the tapering estimator $\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)$, suppose $\{\mathbf{x}_{ti}^{f}\}_{i=1}^{n}$ are sub-Gaussian, then for $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}(\varepsilon_{0},\alpha,C)$, $\mathbb{E}\|\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}.$
@@
    For the thresholding estimator $T_{s}({\mathbf{S}}_{t}^{f})$, for sufficiently large $M^{\prime}$, if $s=M^{\prime}(n^{-1}{\log p})^{-1/2}$, then, uniformly on $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M\right)$, $\mathbb{E}\|T_{s}({\mathbf{S}}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp c_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}.$
@@
According to Proposition 3.1, the convergence results in Proposition 3.2 on the high dimensional estimation of $\boldsymbol{\Sigma}_{t}^{f}$ translate to that of the estimated Kalman Gain $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}(\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t})^{-1}$ to $\mathbf{K}_{t}$. Specifically,
[…]
As will be shown in the following, the estimation error of the Kalman gain matrix $\Delta_{\mathbf{K},t}=\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ plays a crucial role in quantifying the effects of the high dimensional  banding, the mid-banding, the tapering and the thresholding estimators for $\boldsymbol{\Sigma}_{t}^{f}$.
@@ 3.2 Convergence of High Dimensional Ensemble Kalman Filter
When the dimension $p$ of $\mathbf{x}_{t}$ is high, to allow consistency of the data assimilation, we have to use the high dimensional covariance estimators in Section 3.1 for the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ and the Kalman Gain matrix $\mathbf{K}_{t}$. Based on the consistent estimators of $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{K}_{t}$, we propose the HD-EnKF algorithms given in Section S.1 of the SM and we establish the accuracy of the assimilated state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$. To simplify notation, we denote $\Delta_{\mathbf{K},k}=\mathbb{E}\|\hat{\mathbf{K}}_{k}-\mathbf{K}_{k}\|^{2}$ and
@@
$-2\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}/n\big{)}+2\operatorname{tr}\big{(}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}\big{)}.$ (3.2)
The following Theorem 3.1 presents the upper bounds of the discrepancy between the EnKF and the Kalman Filters for the one-step data assimilation.
@@
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f})|\mathbf{y}_{1:t-1}\|^{2}$ $=$ ${n}^{-1}p^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})=O\big{(}n^{-1}\big{)},$
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $=$ $\operatorname{tr}\big{(}n^{-1}p^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+p^{-1}\Omega_{t}$
$=$ $O\big{(}n^{-1}+p^{-1}\Delta_{\mathbf{K},t}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}\big{)}.$
@@
Theorem 3.1 gives the upper bounds of the one-step data assimilation MSE with high dimensional forecast error covariance estimators, which depends on $\Delta_{\mathbf{K},t}$ whose convergence rates for different covariance estimators are given in (3.1). Specifically, with the banding or mid-banding estimators, the tapering estimator and the thresholding estimator, the one-step EnKF data assimilation MSE are, respectively,
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\{n^{-1},\;p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\},$ (3.4)
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;p^{-1}q(n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p})\big{\}},$ (3.5)
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}$ $\asymp\max\big{\{}n^{-1},\;p^{-1}qc_{0}^{2}(p)\left(n^{-1}{\log p}\right)^{(1-\gamma)}\big{\}}.$ (3.6)
@@
For fixed $p,q$, both the forecast and analysis states of the EnKF converge to the underlying forecast and analysis states of the Kalman Filter with a convergence rate of $\sqrt{n}$ as $n\rightarrow\infty$ as $\Delta_{\mathbf{K},t}=O(n^{-1})$, as also shown in Theorem S.3.1 of the SM. Furthermore, Proposition S.3.1 implies that with linear model operator $\mathcal{M}_{t}$ and Gaussian initial state, the asymptotic distributions of $\mathbf{x}_{t,j}^{a}|\mathbf{y}_{1:t}$ and $\mathbf{x}_{t,j}^{f}|\mathbf{y}_{1:t-1}$ in the EnKF are the same as those of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $\mathbf{x}_{t}|\mathbf{y}_{1:t-1}$ in the Kalman Filter, respectively.
@@
To show the impact of the dimensionality on the MSE, we take the banding and mid-banding estimators as examples. If $q\asymp p$, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O((n^{-1}{\log p})^{2\alpha/(2(\alpha+1))})$, which implies that the one-step assimilation MSE mainly depends on the estimation error of the Kalman gain matrix. Commonly in high resolution data assimilation, the dimension of the observation $\mathbf{y}_{t}$ is much smaller than that of the state vector $\mathbf{x}_{t}$, so that $q\ll p$. If $q$ is relatively large such that $p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\}\gg n^{-1}$, then the estimation error of the Kalman gain matrix will dominate, that is, $p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}=O\big{(}p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\big{)}$. Otherwise, the first term $n^{-1}$ dominates. For the impact of the observation dimension $q$ on the MSE, a study reported in the SM shows that if the additional observation $y_{t,q+1}$ is not far away from its forecast ${\mathbf{H}}_{t}\mathbf{x}_{t}^{f}$, the MSE will decrease as the first term in (3.1) will decrease as $q$ increases.
@@
For the discrepancies between the EnKF analysis state and the true state, as $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a})|\mathbf{y}_{1:t}\|^{2}+\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}$ and $\mathbf{x}_{t}$ is a random variable with $\mathbb{E}\|(\mathbf{x}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{a})$ which is a constant not related to the ensemble size $n$ for a fixed $p$, then $\mathbb{E}\|(\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t})|\mathbf{y}_{1:t}\|^{2}$ can not converge to zero.
@@
In practice, the data assimilation process can be repeated for multiple time steps ahead to make longer-time predictions. The following Theorem 3.2 gives the $h$-step MSE bounds between the forecast/analysis states of the EnKF in comparison to the underlying Kalman Filters. To simplify writing, define two non-stochastic constants
@@
where $A_{t,h}=\max_{t<k\leq t+h}\|\mathbf{I}-\mathbf{K}_{k}\mathbf{H}_{k}\|$, $B_{t,h}=\max_{t<k\leq t+h}\{\mathbb{E}\|\mathbf{I}-\hat{\mathbf{K}}_{k}\mathbf{H}_{k}\|\}$, and $L_{t,h}=\max_{0<k\leq h}\{\ell_{0},\ell_{t+k}\}$ for a fixed $t$, with $\ell_{0}$ and $\ell_{t}$ defined in Assumption 1. We further denote
$U_{k}=\big{\{}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\Delta_{\mathbf{K},k}+\operatorname{tr}\big{(}\mathbb{E}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})^{\mathrm{T}}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})({\mathbf{H}}_{k}\boldsymbol{\Sigma}_{k}^{f}{\mathbf{H}}_{k}^{\mathrm{T}}+\mathbf{R}_{k})\big{)}\big{\}}^{1/2}.$ (3.8)
@@
$p^{-1}\mathbb{E}\|(\hat{\mathbf{x}}_{t+h}^{f}-\mathbf{x}_{t+h}^{f})|\mathbf{y}_{1:t+h-1}\|^{2}\leq p^{-1}\Big{(}\{\operatorname{tr}(n^{-1}\boldsymbol{\Sigma}_{t+h}^{f})\}^{1/2}+L_{t,h}C_{t,h-1}^{B}\max_{t<k\leq t+h-1}U_{k}\Big{)}^{2}$

### v3

@@ 3.1 Consistent estimation of high dimensional forecast error covariance matrix
    For the mid-banding estimator $B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})$ with $k_{1}+k_{2}\asymp(n^{-1}\log p)^{-1/(2\alpha+2)}$, uniformly on $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right)$, $\mathbb{E}\|B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp(n^{-1}{\log p})^{\alpha/(\alpha+1)}.$
@@
    For the tapering estimator $\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)$ and $\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}(\varepsilon_{0},\alpha,C)$, the optimal rate of convergence of the tapering estimator is $\mathbb{E}\|\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k)-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp\min\big{\{}n^{-2\alpha/(2\alpha+1)}+{n}^{-1}{\log p},\;{n}^{-1}{p}\big{\}}.$
@@
    For the thresholding estimator $T_{s}({\mathbf{S}}_{t}^{f})$, for sufficiently large $M^{\prime}$, if $s=M^{\prime}(n^{-1}{\log p})^{-1/2}$, then, uniformly on $\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M\right)$, $\mathbb{E}\|T_{s}({\mathbf{S}}_{t}^{f})-\boldsymbol{\Sigma}_{t}^{f}\|^{2}\asymp c_{0}^{2}(p)(n^{-1}{\log p})^{(1-\gamma)}.$
@@
According to Lemma 3.1, the convergence results in Proposition 3.1 on the high dimensional estimation of $\boldsymbol{\Sigma}_{t}^{f}$ translate to that of the estimated Kalman Gain $\hat{\mathbf{K}}_{t}=\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}(\mathbf{H}_{t}\hat{\boldsymbol{\Sigma}}_{t}^{f}\mathbf{H}_{t}^{\mathrm{T}}+\mathbf{R}_{t})^{-1}$ to $\mathbf{K}_{t}$. Specifically,
$\text{ if }\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right)\text{or }\mathcal{U}_{bc}\left(\varepsilon_{0},\alpha,C\right),\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}\asymp(\frac{\log p}{n})^{\frac{\alpha}{\alpha+1}}\,\text{ for }\hat{\boldsymbol{\Sigma}}_{t}^{f}=B_{k}(\mathbf{S}_{t}^{f})\text{ or }B_{k_{1},k_{2}}(\mathbf{S}_{t}^{f});$
$\text{ if }\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{b}\left(\varepsilon_{0},\alpha,C\right),\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}\asymp\min\big{\{}n^{\frac{-2\alpha}{2\alpha+1}}+\frac{\log p}{n},\;\frac{p}{n}\big{\}}\text{ for }\,\hat{\boldsymbol{\Sigma}}_{t}^{f}=\hat{\boldsymbol{\Sigma}}_{t,tap}^{f}(k);$
$\text{ if }\boldsymbol{\Sigma}_{t}^{f}\in\mathcal{U}_{\tau}\left(\gamma,c_{0}(p),M\right),\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}\asymp c_{0}^{2}(p)(\frac{\log p}{n})^{(1-\gamma)}\ \text{ for }\hat{\boldsymbol{\Sigma}}_{t}^{f}=T_{s}(\mathbf{S}_{t}^{f}).$ (3.1)
As will be shown later, the Kalman Gain estimation error $\Delta_{\mathbf{K},t}=\mathbb{E}\|\hat{\mathbf{K}}_{t}-\mathbf{K}_{t}\|^{2}$ plays a crucial role in quantifying the effects of the high dimensional  banding, mid-banding, tapering and thresholding estimators for $\boldsymbol{\Sigma}_{t}^{f}$.
@@ 3.2 Convergence of High Dimensional Ensemble Kalman Filter
When the dimension $p$ of $\mathbf{x}_{t}$ is high, to allow consistency of the data assimilation, we have to use the high dimensional covariance estimators in Section 3.1 for the forecast error covariance matrix $\boldsymbol{\Sigma}_{t}^{f}$ and the Kalman Gain matrix $\mathbf{K}_{t}$. Based on the consistent estimators of $\boldsymbol{\Sigma}_{t}^{f}$ and $\mathbf{K}_{t}$, we propose the HD-EnKF algorithms given in Section S.1 of the SM and we establish the accuracy of the assimilated state $\hat{\mathbf{x}}_{t}^{a}$ to the underlying $\mathbf{x}_{t}^{a}$ in the following. To simplify notation, we denote $\Delta_{\mathbf{K},k}=\mathbb{E}\|\hat{\mathbf{K}}_{k}-\mathbf{K}_{k}\|^{2}$ and
@@
$-2\operatorname{tr}\big{(}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{H}_{t}\boldsymbol{\Sigma}_{t}^{f}/n\big{)}+2\operatorname{tr}\big{(}\big{(}\mathbf{K}_{t}^{\mathrm{T}}\mathbb{E}(\hat{\mathbf{K}}_{t}-\mathbf{K}_{t})\mathbf{R}_{t}/n\big{)}\big{)}.$
The following Theorem 3.1 presents the upper bounds of the discrepancy between the EnKF analysis state and the underlying Kalman Filter analysis for the one-step data assimilation for a general case where the model operator $\mathcal{M}_{t}$ is possibly nonlinear.
@@
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{f}-\mathbf{x}_{t}^{f}\|^{2}|\mathbf{y}_{1:t-1}\}={n}^{-1}p^{-1}\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{f})=O\big{(}n^{-1}\big{)},$
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=\operatorname{tr}\big{(}n^{-1}p^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}\big{)}+p^{-1}\Omega_{t}$ (3.2)
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\;\,=O\big{(}n^{-1}+p^{-1}\Delta_{\mathbf{K},t}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2}\big{)}.$
@@
Theorem 3.1 gives the upper bounds of the one-step data assimilation MSE with high dimensional forecast error covariance estimators, which depends on $\Delta_{\mathbf{K},t}$ whose convergence rates for different covariance estimators are given in (3.1). Note that the first term in (3.2), $n^{-1}(\mathbf{I}-\mathbf{K}_{t}\mathbf{H}_{t})\boldsymbol{\Sigma}_{t}^{f}=n^{-1}\boldsymbol{\Sigma}_{t}^{a}$, is the variance of the sample mean of the analysis ensemble using the true forecast error covariance $\boldsymbol{\Sigma}_{t}^{f}$ and the second term $\Omega_{t}$ results from the estimation error of $\hat{\boldsymbol{\Sigma}}_{t}^{f}$. Specifically, with the banding or mid-banding estimators, the tapering estimator and the thresholding estimator, the one-step EnKF data assimilation MSE are, respectively,
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\{n^{-1},\;p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\},$ (3.3)
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\big{\{}n^{-1},\;p^{-1}q(n^{\frac{-2\alpha}{2\alpha+1}}+{n}^{-1}{\log p})\big{\}},$ (3.4)
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}$ $\asymp\max\big{\{}n^{-1},\;p^{-1}qc_{0}^{2}(p)(n^{-1}{\log p})^{(1-\gamma)}\big{\}}.$ (3.5)
It should also be noted that the above results are valid for general model operators $\mathcal{M}_{t}$ given in Assumption 1-(i) and (ii). Compared with the results in Al-Ghattas and Sanz-Alonso (2024), they investigated the difference between the ensembles of single-step Ensemble Kalman Inversion using the localized estimation and the true forecast error covariance matrices, under the framework of the effective dimension of a matrix $\mathbf{A}$ defined by $\operatorname{tr}(\mathbf{A})/\|\mathbf{A}\|$ (or the max-log effective dimension for a sparse covariance).
@@
For fixed $p,q$, both the forecast and analysis states of the EnKF converge to the underlying forecast and analysis states of the Kalman Filter with a convergence rate of $\sqrt{n}$ as $n\rightarrow\infty$ because $\Delta_{\mathbf{K},t}=O(n^{-1})$, as also shown in Theorem S.3.1 of the SM. Furthermore, for linear model operator ${\mathbf{M}}_{t}$, Proposition S.3.1 of the SM implies that with Gaussian initial state, the asymptotic conditional distributions of the analysis ensemble $\mathbf{x}_{t,j}^{a}|\mathbf{y}_{1:t}$ and the forecast ensemble $\mathbf{x}_{t,j}^{f}|\mathbf{y}_{1:t-1}$ in the EnKF are the same as those of $\mathbf{x}_{t}|\mathbf{y}_{1:t}$ and $\mathbf{x}_{t}|\mathbf{y}_{1:t-1}$ in the Kalman Filter, respectively.
@@
To show the impact of the dimensionality on the assimilation MSE, we take the banding and mid-banding estimators as examples. If $q\asymp p$, $p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O((n^{-1}{\log p})^{2\alpha/(2\alpha+2)})$, which implies that the one-step assimilation MSE mainly depends on the estimation error of the Kalman Gain. Commonly in high resolution data assimilation, the dimension of the observation $\mathbf{y}_{t}$ is much smaller than that of the state vector $\mathbf{x}_{t}$, so that $q\ll p$. If $q$ is relatively large such that $p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\}\gg n^{-1}$, then the estimation error of the Kalman Gain matrix will dominate, that is, $p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O\big{(}p^{-1}q(n^{-1}{\log p})^{\alpha/(\alpha+1)}\big{)}$. Otherwise, the first term $n^{-1}$ dominates. However, if sample covariance $\mathbf{S}_{t}^{f}$ is used in $\hat{\mathbf{K}}_{t}$ when $p\geq n$, then $p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}=O(p/n^{2}\|\mathbf{y}_{t}-{\mathbf{H}}_{t}\mathbf{x}_{t}^{f}\|^{2})=O(pq/n^{2})$. For the impact of the observation dimension $q$ on the MSE, a study in the SM shows that if the additional observation $y_{t,q+1}$ is not far away from its forecast ${\mathbf{H}}_{t}\mathbf{x}_{t}^{f}$, the MSE will decrease since the first term in (3.2) decreases more.
@@
For the discrepancies between the EnKF analysis state and the true state, as $\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}\|^{2}|\mathbf{y}_{1:t}\}=\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}^{a}\|^{2}|\mathbf{y}_{1:t}\}+\mathbb{E}\{\|\mathbf{x}_{t}^{a}-\mathbf{x}_{t}\|^{2}|\mathbf{y}_{1:t}\}$ and $\mathbf{x}_{t}$ is a random variable with $\mathbb{E}\{\|\mathbf{x}_{t}^{a}-\mathbf{x}_{t}\|^{2}|\mathbf{y}_{1:t}\}=\operatorname{tr}(\boldsymbol{\Sigma}_{t}^{a})$ which is a constant not related to the ensemble size $n$ for a fixed $p$, then $\mathbb{E}\{\|\hat{\mathbf{x}}_{t}^{a}-\mathbf{x}_{t}\|^{2}|\mathbf{y}_{1:t}\}$ can not converge to zero. This is due to the fact that the underlying state is a random variable. In contrast, $\mathbf{x}_{t}^{a}$ is a conditional mean and is possible to be estimated consistently.
@@
In practice, the ensemble forecast-analysis process in the EnKF repeats for multiple time steps ahead to make longer-time predictions. Thus, in addition to the one-step ahead data assimilation accuracy, one may also care about the quality of multiple-step forecasting and analysis. The following Theorem 3.2 gives the $h$-step MSE bounds between the forecast/analysis states of the EnKF in comparison to the underlying Kalman Filters. To simplify writing, define two non-stochastic constants
@@
where $A_{t,h}=\max_{t<k\leq t+h}\|\mathbf{I}-\mathbf{K}_{k}\mathbf{H}_{k}\|$, $B_{t,h}=\max_{t<k\leq t+h}\{\mathbb{E}\|\mathbf{I}-\hat{\mathbf{K}}_{k}\mathbf{H}_{k}\|\}$, and $L_{t,h}=\max_{0<k\leq t+h}\{\ell_{0}$,
$\ell_{k}\}$ with $\ell_{0}$ and $\ell_{k}$ being respectively the Lipschitz coefficient and the upper bound of the local linearization coefficient of the model operation defined in Assumption 1. We further denote
$U_{k}=\big{\{}\|\mathbf{y}_{k}-{\mathbf{H}}_{k}\mathbf{x}_{k}^{f}\|^{2}\Delta_{\mathbf{K},k}+\operatorname{tr}\big{(}\mathbb{E}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})^{\mathrm{T}}(\hat{\mathbf{K}}_{k}-\mathbf{K}_{k})({\mathbf{H}}_{k}\boldsymbol{\Sigma}_{k}^{f}{\mathbf{H}}_{k}^{\mathrm{T}}+\mathbf{R}_{k})\big{)}\big{\}}^{{1}/{2}}.$
@@
$p^{-1}\mathbb{E}\{\|\hat{\mathbf{x}}_{t+h}^{f}-\mathbf{x}_{t+h}^{f}\|^{2}|\mathbf{y}_{1:t+h-1}\}\leq p^{-1}\Big{(}\{\operatorname{tr}(n^{-1}\boldsymbol{\Sigma}_{t+h}^{f})\}^{1/2}+L_{t,h}C_{t,h-1}^{B}\max_{t<k\leq t+h-1}U_{k}\Big{)}^{2}$